OpenStudy (fairy733189):

The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n) = -250n2 + 2,500n - 5,250 Part A: What are the zeroes of the above function, and what do they represent? Show your work. Part B: Find the maximum profit by completing the square of the function P(n). Show the steps of your work. Part C: What is the axis of symmetry of the function P(n)?

OpenStudy (fairy733189):

@Mehek14

satellite73 (satellite73):

\[P(n) = -250n^2 + 2,500^n - 5,250 \] factor out the \(250\) first, that will make finding the zeros easier

satellite73 (satellite73):

actually even better is to factor out the \(-250\)

OpenStudy (fairy733189):

@satellite73

OpenStudy (fairy733189):

would it be -250(n^2 + 10 - 21)?

OpenStudy (fairy733189):

@mathmale

OpenStudy (fairy733189):

Is the correct answer -250(n^2 - 10 +21)?

OpenStudy (mathmale):

"Part A: What are the zeroes of the above function, and what do they represent? Show your work." This involves setting the given profit function = to 0 and solving for the zeros / roots / solutions. First, however, I'd like for you to verify that you've copied down this problem correctly:\[P(n) = -250n2 + 2,500n - 5,250\]

OpenStudy (fairy733189):

ugh sooorrry I alays forget to put the ^ sumbol its supposed to be -250n^2 + 2500n - 5250

OpenStudy (mathmale):

should be written as P(n) = -250n^2 + 2,500n - 5,250. This is a quadratic that's pretty straightforward to solve for roots. Somehow your P(n) = -250n2 + 2,500n - 5,250 became \[P(n) = -250n^2 + 2,500^n - 5,250,\]

OpenStudy (mathmale):

which is dramatically different from a quadratic because you're using n as an exponent. Try again. What is the equation after you've factored out the 250?

OpenStudy (fairy733189):

would it be -250(n^2 - 10 + 21) -250(n -7)(n-3)

OpenStudy (mathmale):

You need to find the roots / zeros / solutions of P(n) = -250n2 + 2,500n - 5,250 = 250(-n^2 + 10n - 21) = 0.

OpenStudy (mathmale):

Be certain you know why your 10 is not correct, but 10n would be.

OpenStudy (fairy733189):

right sorry. So do I factor completely or do I factor the 250 and then make the new equation equal 0?

OpenStudy (mathmale):

Yes. Factor out -250, actually. Once you've done that, set the resulting expression = to zero and factor it (or use some other method to solve this quadratic equation. P(n)=-250( ???? ) = 0

OpenStudy (fairy733189):

it would be p(n) = -250(n^2 - 10n + 21)

OpenStudy (fairy733189):

and = 0 would come after

OpenStudy (fairy733189):

but how do I factor that?

OpenStudy (mathmale):

Focus on your n^2 - 10n + 21 = 0. Look at the last term; it's 21. What are possible factors, either + or -, of 21? Hopefully your factors, when added together, will sum up to -10n.

OpenStudy (fairy733189):

the two factors would be -7 and -3

OpenStudy (mathmale):

Yes. Multiply them together, and you get +21 (which is the 3rd term of your quadratic). combine them and you get -10n, which is the 2nd term of your quadratic. So far, so good!

OpenStudy (mathmale):

Actually, "the two factors would be -7 and -3" is not quite right; you meant: "The two roots would be -7 and -3. If -7 is a root, a factor of that quadratic would be (x+7).

OpenStudy (fairy733189):

then would the new equation be (n - 3)(n - 7)?

satellite73 (satellite73):

yes

satellite73 (satellite73):

or rather \[(n-3)(n-7)=0\] which you can solve in your head for \(n\)

OpenStudy (mathmale):

Fairy: Earlier, you typed "the two factors would be -7 and -3" I should have caught that. The roots are 7 and 3, and the factors are (x-7) and (x-3). To verify this, multiply those two factors together.

OpenStudy (mathmale):

How would you now find the axis of symmetry of this parabolic graph? Hint: have you seen this formula before? x=-b/(2a)

OpenStudy (fairy733189):

Yes but is this still for part a?

OpenStudy (fairy733189):

would the answer to part a be the zeroes are -7 and -3 or 7 and 3?

OpenStudy (fairy733189):

@mathmale

OpenStudy (fairy733189):

I think it would be 7 and 3 right since money can't be negative?

satellite73 (satellite73):

yes good point

OpenStudy (fairy733189):

so 7 and 3 are the two zeroes?

satellite73 (satellite73):

yes but be careful

satellite73 (satellite73):

Part A: What are the zeroes of the above function, and what do they represent? Show your work.

satellite73 (satellite73):

they make no money in profit if they sell the tickets for 3 or 7 dollars

OpenStudy (fairy733189):

what do you mean?

OpenStudy (fairy733189):

you make a good point but part a is asking for the zeroes of the function

satellite73 (satellite73):

The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each:

satellite73 (satellite73):

if they sell for \(3\) then \(P(3)=0\) they make no profit if they sell for \(7\) then \(P(7)=0\) still no profit

OpenStudy (fairy733189):

oh ok thanks! can you help me with part b?

satellite73 (satellite73):

sure

satellite73 (satellite73):

first off the answer is obvious the function is a quadratic that opens down it is zero at 3 and 7, so it has a maximum half way between them what is half way between 3 and 7?

satellite73 (satellite73):